OpenStudy (anonymous):

how do you solve equations like this? 2x^2=32

jimthompson5910 (jim_thompson5910):

hint: divide both sides by 2 to get 2x^2=32 2x^2/2=32/2 1x^2 = 16 x^2 = 16 x = ??? or x = ???

OpenStudy (anonymous):

^^^^ what you have is good so far, now just take the square root of x^2 and 16 so that it just becomes: x = 4

OpenStudy (anonymous):

ohhhhh okay omg thank you!

OpenStudy (anonymous):

can we try another one real quick and see if i got this down?

jimthompson5910 (jim_thompson5910):

remember that if x^2 = 16, then x = 4 or x = -4

jimthompson5910 (jim_thompson5910):

there are two solutions here

jimthompson5910 (jim_thompson5910):

this is because 4^2 = 16 and (-4)^2 = 16

OpenStudy (anonymous):